# Piezoelectric actuation for integrated photonics

### v1

@@ 3.3.2 Unreleased PZT Tuning of Si3N4 Microring Resonator
The PZT tuning of Si3N4 photonic devices has been explored in the past decade, as the fabrication of high-quality, low-loss Si3N4 waveguide matures [224, 213, 214, 24, 223]. Two kinds of Si3N4 waveguide design are widely explored. One is the thick Si3N4 waveguide which has tight mode confinement and is fabricated by the photonic Damascene process [224, 213, 215, 24] to address the high stress built in thick Si3N4 film. The other design uses a thin Si3N4 waveguide fabricated by the subtractive process where the optical mode mainly resides in the SiO2 cladding [211, 210, 5]. Since the mode is less confined, the radius of the thin Si3N4 optical ring resonator is on the order of millimeter for maintaining low bending loss, while the radius of thick Si3N4 ring resonator can be as low as 20 m [227].
@@
The tuning efficiency increases as the device’s size decreases (FSR increases). This is due to the contribution from the relative change of the radius to the resonant frequency (see Eq.13), which is larger for smaller radius . For the 25-GHz-FSR device (red curve), over half FSR tuning is obtained under 40 V, leading to phase shift of the light in the microring. The figure of merit (modulation length product) is calculated to be 22 Vcm, which is an order of magnitude larger than the state of the electro-optic modulation using LiNbO3 [10]. However, the modulation loss product VdB is comparable with LiNbO3 (1 VdB), thanks to the ultra-low loss of the Si3N4 waveguide. The tuning speed can be inferred from the time response as shown in Fig. 19(c-d). For a smaller device (FSR = 100 GHz), it takes 0.2 s to switch from one state to the other. Due to the large permittivity of PZT, the speed is ultimately limited by the large RC time constant for the large devices. Thus, the 30 GHz FSR device takes 1.4 s to respond, which is the time to charge the large PZT capacitor.
@@
The piezoelectric tuning of thin Si3N4 waveguide (175 nm thick) is demonstrated by Wang et al. [107] as shown in Fig. 20. Due to the low confinement of the thin Si3N4 waveguide, to maintain the high intrinsic optical , a gap is opened in the PZT actuator right above the waveguide, and the PZT is offset by 2 m relative to the center of the waveguide [see Fig. 20(a)]. The optical image of the fabricated device with 625 m radius is shown in Fig. 20(b). The DC tuning of the resonant wavelength under forward and backward voltage sweep is plotted in Fig. 20(c). A nonlinear and hysteresis tuning can be observed with tuning efficiency of 162 MHz/V, which leads a of 43 Vcm and of 1.3 VdB. Similarly, a 20 nW power consumption at 20 V was reported.
@@
Interestingly, a butterfly-like curve of the tuning-voltage is obtained which agrees with the typical strain-electric field curve for ferroelectric materials. It is related with the switching of polarization domains under the electric field. For example, as the voltage increases from negative to zero (blue curve), the population of domains that align in the negative direction gradually decreases but still with some remaining at . As the voltage further changes to positive sign but at small value, the electric field and the polarization have the opposite sign and the wavelength increases (relative to ). As the voltage further increases, all the domains align with the electric field and now in the positive direction, and then the wavelength changes from increasing to decreasing. The coercive field of the PZT can be found at the turning point where all the domains are switched by the field. Similar effect can be observed when sweeping the voltage backward. Therefore, different from the AlN, the PZT shows the same tuning direction at high positive and negative voltages since the domains are kept aligned with the field. To achieved bi-directional tuning, a DC bias needs to be applied and maintained. 3-dB tuning bandwidth of 15 MHz (corresponding to 0.067 s) is achieved, due to the donut-shaped actuator which has smaller area than the disk. Therefore, depending on the specific application, the shape and area of the actuator designed carefully, trading-off between tuning efficiency and speed.
@@ 3.3.3 Unreleased PZT Tuning of Si3N4 Mach-Zehnder Interferometer
Besides the microring resonator, the optical Mach-Zehnder interferometer (MZI) has been another important building block in integrated optical circuits. Different from the microring which can only operate at discrete optical resonant frequencies with narrow linewidth, the MZI can support a wide optical bandwidth at the expense of larger footprint. This feature helps it to stand out in electro-optic modulators which usually shows higher modulation bandwidth than the microring [10] and is easier to implement wavelength-division multiplexing (WDM) in optical telecommunication. Moreover, MZI can conduct beam-splitter operation on chip, which is the basic element in optical neural networks [55] and photonic quantum computing [56, 229]. Active tuning of the MZI is highly desired for programmable photonic circuits to be reconfigured to conduct complex computing tasks, such as unitary transformations [230, 231, 232]. However, most of the demonstrations rely on thermo-optic tuning, which is practically limited in scaling, power consumption, and cross-talk [55, 56, 232]. In that sense, piezoelectric tuning is gaining much attention and will play an important role in the future.
The tuning of Si3N4 Mach-Zehnder interferometer based on PZT actuators was pioneered by Heideman et al. [91] as shown in Fig. 21(a), where a rectangular PZT actuator is placed on top of one arm of the MZI. The stress induced by the PZT will change the effective refractive index of the waveguide and thus the relative phases between the two arms and the output light intensity after interference. A 25 nm thin Si3N4 waveguide (4 m wide) is designed to support the fundamental TM mode at an optical wavelength of 640 nm. Due to the weak confinement of the optical mode, the waveguide is placed 8 m below the bottom metal to maintain a low optical loss. Unlike the works shown in previous sections, 2 m PZT is deposited by pulsed-laser deposition (PLD). Note that the PZT uniformly covers the whole device, and only the top metal is patterned to generate the local stress.
@@
The phase shift of the MZI was further improved by the same group recently by optimizing the design of both the PZT actuator and the Si3N4 waveguide [233, 90]. As shown in Fig. 21(d), a dome-shape actuator was introduced which can concentrate the stress efficiently on the waveguide as can be seen from the numerical simulation in Fig. 21(e). This result agrees with the previous finding of the work on PZT tuning of the Si microring resonator [106] (see Fig. 18). Moreover, the dome can be naturally formed during the subtractive fabrication of the Si3N4 waveguide, where etching the Si3N4 leaves a surface topology with a height of 1 m and the dome is formed by uniformly depositing SiO2 top cladding covering the whole waveguide. On the other hand, an asymmetric double-strip waveguide was designed to tightly confine the optical mode so that the thickness of the top cladding was reduced from 8 m to 3 m. This can largely increase the overlap between the optical mode and the stress distribution. With these improvements being implemented, the phase shift between 1 to 2 can be achieved under 40 V, depending on the structural properties of the deposited PZT films, as shown in Fig. 21(f). The maximum of 16 V cm is nearly twice better than the PZT actuator with flat structure. The tuning speed was further inferred from the time domain step response with a fitted time constant of 0.85 s. The high speed benefits from the dome-shape design which allows narrower actuator (10 m in this case) with smaller area and capacitance [233].
@@ 3.4 Comparison of the State-of-the-Arts
In this section, we summarize and compare the performance of piezoelectric tuning of integrated photonic devices demonstrated with AlN and PZT actuators, providing a guidance for future device design, as shown in Table 3. It can be seen PZT actuators generally shows an order of magnitude higher tuning efficiency than AlN, because of its larger piezoelectric response. However, AlN actuator shows much more linear response, negligible hysteresis, low leakage current, bi-directional tuning (which means zero biasing), and compatibility with CMOS processes. These features are highly desired in applications that require fast and repeatable transduction of input signals, such as frequency-modulated continuous-wave (FMCW) LiDAR [38, 52]. On the other hand, released actuator has larger tuning range than unreleased device, but at the expense of narrower tuning bandwidth. The currently most efficient tuning is achieved by combing the PZT and releasing process, as demonstrated by Ref. [88]. Notably, Si waveguide shows several times larger tuning than Si3N4 because of its higher photoelastic coefficient and refractive index. However, Si waveguide demonstrated thus far has propagation loss much higher than the Si3N4. With further optimization of the fabrication to suppress the losses, piezoelectric tunable Si photonics could become a promising platform for programmable photonic networks [221, 229, 55] and reconfigurable optical signal processors.
@@ 4 Piezoelectric Tuning in Laser Applications
Due to its low optical losses and high Kerr nonlinearity, Si3N4 has become one of the main material platforms in hybrid integrated lasers with narrow linewidth [5] and soliton microcombs with low pump threshold [49]. While modulators integrated with Si3N4, e.g. based on the electro-absorption of graphene [200, 71, 70, 202] and ferroelectric lead-zirconate-titanate (PZT) [75] have been demonstrated, they have not shown compatibility with soliton generation which requires maintaining ultralow optical losses. As a result, current integrated actuation techniques for soliton microcombs rely primarily on metallic heaters [63, 62], which have limited actuation bandwidth up to 10 kHz. Furthermore, heaters exhibit unidirectional tuning and typically consume electrical power exceeding 30 mW [39], higher than the optical power threshold for soliton formation in state-of-the-art integrated devices [39, 40], and are not compatible with cryogenic operation [65, 241]. Alternatively, the direct generation of solitons in Pockels materials such as AlN [216, 46, 47, 242] and LiNbO3 [45, 44, 43] can allow high-speed actuation, however these platforms are not yet as mature as Si3N4.
@@ 4.1 Narrow-Linewidth, Frequency-Agile Integrated Laser
Over the past decades, the heterogeneous integrated semiconductor laser has revolutionized the application of laser in our daily life, from a laser pointer to a data center, due to its compact size and low cost [243]. However, its frequency noise is still higher than the fiber laser. It is only most recently that the coupling to a low-loss Si3N4 photonic circuit has made it possible for a chip-scale, low-noise laser, thanks to the maturity of hybrid and heterogeneous integration [244]. A variety of coupling schemes has been invented, such as wafer bonding [7, 4, 245, 246], edge coupling [41, 5, 247], and photonic wire bonding [248, 249, 250]. A Hertz-linewidth laser has been achieved by properly tuning the laser injection current such that the CW laser’s frequency coincides with the resonance of an external Si3N4 microresonator [5, 251]. The laser emission frequency is then self-injection-locked to the Si3N4 microresonator, a phenomenon in which the laser dynamics is impacted by external feedback light in the case without an optical isolator [252]. When light is coupled into the microresonator, bulk and surface Rayleigh scattering inside the microresonator can backscatter some portion of the light and inject it into the laser, which triggers laser self-injection locking and impacts the laser dynamics. The consequence is that the new laser emission frequency is tightly locked to the microresonator resonance whose frequency is close to the initial free-running laser’s emission frequency. The laser linewidth and noise are significantly reduced in the presence of this passive optical feedback (in contrast to active locking of laser frequency such as the PDH lock), and the linewidth is ultimately limited by the thermal-refractive noise (TRN) of the Si3N4 microresonator [253].
@@ 4.1.1 Piezoelectric Tunable Self-Injection Locked DFB Laser
Figure 23(d) shows the laser frequency tuning under a triangularly modulated voltage (150 V) applied on the AlN actuator. The time-varying laser frequency is generally characterised by measuring a heterodyne beatnote with a reference ECDL on a fast photodetector, which produces a time-frequency spectrogram. Consequently, the laser output frequency is triangularly chirped with frequency excursion of 1.1 GHz and speed up to 0.8 MHz. Due to the linearity of the AlN actuator, the frequency chirping nonlinearity is below 10 MHz (defined as the Root Mean Square (RMS) deviation of the measured frequency from an ideal triangular wave calculated by least-squares fitting), and no tuning hysteresis is observed during continuous operation. The degradation of the deviation is primarily due to the nonlinearity of the voltage amplifier at high frequency. This highly linear tuning behavior facilitates the generation of chirps without the need for active or passive linearization, which is ideal for direct implementation in long-range FMCW LiDAR systems [256, 257, 38]. As discussed in the previous chapter, the modulation speed is limited by the chip’s mechanical modes, such as the contour modes [258] and HBAR modes [12]. In Ref. [52], a flat modulation response is extended to 10 MHz by implementing a differential drive on an apodized chip with carbon tape. One should also note that the ultimate speed is limited by the speed of the optical feedback for laser SIL.
In addition to AlN, other piezoelectric materials with larger piezoelectric coefficient, such as lead zirconium titanate (PZT), can be used to reduce the required voltage for a given frequency excursion. The same PZT actuator shown in Fig.19 was used by the same group to tune the frequency of a SIL DFB laser, as demonstrated in Fig. 24(a,b). A 525 MHz frequency excursion at the repetition rate of 500 kHz is achieved by applying only 1.1 V triangular wave. Interestingly, while it is well known that PZT is ferroelectric and exhibits hysteresis, the measured RMS nonlinearity is 5 MHz, and the relative nonlinearity, defined as the ratio of the nonlinearity to the total frequency excursion, is less than 1. This is due to the small sweeping voltage, which is also compatible with CMOS integrated circuits (e.g., application-specific integrated circuit (ASIC)), paving the way for a fully photonic-electronic integrated tunable hybrid laser systems [260].

### v2

@@ 3.3.2 Unreleased PZT Tuning of Si3N4 Microring Resonator
The PZT tuning of Si3N4 photonic devices has been explored in the past decade, as the fabrication of high-quality, low-loss Si3N4 waveguide matures [235, 224, 225, 24, 234]. Two kinds of Si3N4 waveguide design are widely explored. One is the thick Si3N4 waveguide, which has tight mode confinement and is fabricated by the photonic Damascene process [235, 224, 226, 24] or the thermal cycling process to address the high stress built in thick Si3N4 film [238, 239]. The other design uses a thin Si3N4 waveguide fabricated by the subtractive process where the optical mode mainly resides in the SiO2 cladding [222, 221, 5]. Since the mode is less confined, the radius of the thin Si3N4 optical ring resonator is on the order of millimeter for maintaining low bending loss, while the radius of thick Si3N4 ring resonator can be as low as 20 m [240].
@@
The tuning efficiency increases as the device’s size decreases (FSR increases). This is due to the contribution from the relative change of the radius to the resonant frequency (see Eq.14), which is larger for smaller radius . For the 25-GHz-FSR device (red curve), over half FSR tuning is obtained under 40 V, leading to phase shift of the light in the microring. The figure of merit (modulation length product) is calculated to be 22 Vcm, which is an order of magnitude larger than the state of the electro-optic modulation using LiNbO3 [10]. However, the modulation loss product VdB is comparable with LiNbO3 (1 VdB), thanks to the ultra-low loss of the Si3N4 waveguide. The tuning speed can be inferred from the time response as shown in Fig. 19(c-d). For a smaller device (FSR = 100 GHz), it takes 0.2 s to switch from one state to the other. Due to the large permittivity of PZT, the speed is ultimately limited by the large RC time constant for the large devices. Thus, the 30 GHz FSR device takes 1.4 s to respond, which is the time to charge the large PZT plate capacitor.
@@
The piezoelectric tuning of thin Si3N4 waveguide (175 nm thickness) is demonstrated by Wang et al. [109] as shown in Fig. 20. Due to the low confinement of the thin Si3N4 waveguide, to maintain the high intrinsic optical , a gap is opened in the PZT actuator right above the waveguide, and the PZT is offset by 2 m relative to the center of the waveguide [see Fig. 20(a)]. The optical image of the fabricated device with 625 m radius is shown in Fig. 20(b). The DC tuning of the resonant wavelength under forward and backward voltage sweep is plotted in Fig. 20(c). A nonlinear and hysteresis tuning can be observed with tuning efficiency of 162 MHz/V, which leads a of 43 Vcm and of 1.3 VdB. Similarly, a 20 nW power consumption at 20 V was reported.
@@
Interestingly, a butterfly-like curve of the tuning-voltage is obtained which agrees with the typical strain-electric field curve for ferroelectric materials. It is related with the switching of polarization domains under the electric field. For example, as the voltage increases from negative to zero (blue curve), the population of domains that align in the negative direction gradually decreases but still with some remaining at . As the voltage further changes to positive sign but at small value, the electric field and the polarization have the opposite sign and the wavelength increases (relative to ). As the voltage further increases, all the domains align with the electric field and now in the positive direction, and then the wavelength changes from increasing to decreasing. The coercive field of the PZT can be found at the turning point where all the domains are switched by the field. Similar effect can be observed when sweeping the voltage backward. Therefore, different from the AlN, the PZT shows the same tuning direction at high positive and negative voltages since the domains are kept aligned with the field. To achieved bi-directional tuning, a DC bias needs to be applied and maintained. A 3-dB tuning bandwidth of 15 MHz (corresponding to 0.067 s) is achieved, due to the donut-shaped actuator which has smaller area than the disk. Therefore, depending on the specific application, the shape and area of the actuator should be designed carefully, trading-off between tuning efficiency and speed.
@@ 3.3.3 Unreleased PZT Tuning of Si3N4 Mach-Zehnder Interferometer
Besides the microring resonator, the optical MZI has been another important building block in integrated optical circuits. Different from the microring which can only operate at discrete optical resonant frequencies with narrow linewidth, the MZI can support a wide optical bandwidth at the expense of a larger footprint. This feature helps it to stand out in electro-optic modulators where it usually shows higher modulation bandwidth than the microring [10] and is easier to implement wavelength-division multiplexing (WDM) in optical telecommunication. Moreover, MZI can conduct beam-splitter operation on chip, which is the basic element in optical neural networks [57] and photonic quantum computing [58, 241]. Active tuning of the MZI is highly desired for programmable photonic circuits to be reconfigured to conduct complex computing tasks, such as unitary transformations [242, 243, 244]. However, most of the demonstrations rely on thermo-optic tuning, which is practically limited in scaling, power consumption, and cross-talk [57, 58, 244]. In that sense, piezoelectric tuning is gaining much attention and will play an important role in the future.
The tuning of Si3N4 Mach-Zehnder interferometer based on PZT actuators was pioneered by Heideman et al. [93] as shown in Fig. 21(a), where a rectangular PZT actuator is placed on top of one arm of the MZI. The stress induced by the PZT will change the effective refractive index of the waveguide and thus the relative phases between the two arms and the output light intensity after interference. A 25 nm thin Si3N4 waveguide (4 m wide) is designed to support the fundamental TM00 mode at an optical wavelength of 640 nm. Due to the weak confinement of the optical mode, the waveguide is placed 8 m below the bottom metal to maintain a low optical loss. Unlike the works shown in previous sections, 2 m PZT is deposited by pulsed-laser deposition (PLD). Note that the PZT uniformly covers the whole device, and only the top metal is patterned to generate the local stress.
@@
The phase shift of the MZI was further improved by the same group recently by optimizing the design of both the PZT actuator and the Si3N4 waveguide [245, 92]. As shown in Fig. 21(d), a dome-shape actuator was introduced which can concentrate the stress efficiently on the waveguide as can be seen from the numerical simulation in Fig. 21(e). This result agrees with the previous finding of the work on PZT tuning of the Si microring resonator [108] (see Fig. 18). Moreover, the dome can be naturally formed during the subtractive fabrication of the Si3N4 waveguide, where etching the Si3N4 leaves a surface topology with a height of 1 m and the dome is formed by uniformly depositing SiO2 top cladding covering the whole waveguide. On the other hand, an asymmetric double-strip waveguide was designed to tightly confine the optical mode so that the thickness of the top cladding was reduced from 8 m to 3 m. This can largely increase the overlap between the optical mode and the stress distribution. With these improvements being implemented, the phase shift between 1 to 2 can be achieved under 40 V, depending on the structural properties of the deposited PZT films, as shown in Fig. 21(f). The maximum of 16 V cm is nearly twice better than the PZT actuator with flat structure. The tuning speed was further inferred from the time domain step response with a fitted time constant of 0.85 s. The high speed benefits from the dome-shape design which allows for a narrower actuator (10 m in this case) with smaller area and capacitance [245].
@@ 3.4 Comparison of the State-of-the-Arts
In this section, we summarize and compare the performance of piezoelectric tuning of integrated photonic devices demonstrated with AlN and PZT actuators, providing a guidance for future device design, as shown in Table 3. It can be seen that PZT actuators generally show an order of magnitude higher tuning efficiency than AlN, because of its larger piezoelectric response. However, the AlN actuator shows much more linear response, negligible hysteresis, low leakage current, bi-directional tuning (which means zero biasing), and compatibility with CMOS processes. These features are highly desired in applications that require fast and repeatable transduction of input signals, such as FMCW LiDAR [38, 52]. On the other hand, released actuator has larger tuning range than unreleased device, but at the expense of narrower tuning bandwidth. The currently most efficient tuning is achieved by combing the PZT and releasing process, as demonstrated by Ref. [90]. Notably, Si waveguide shows several times larger tuning than Si3N4 because of its higher photoelastic coefficient and refractive index. However, Si waveguide demonstrated thus far has propagation loss much higher than the Si3N4. With further optimization of the fabrication to suppress the losses, piezoelectric tunable Si photonics could become a promising platform for programmable photonic networks [232, 241, 57] and reconfigurable optical signal processors.
@@ 4 Piezoelectric Tuning in Laser Applications
Due to its low optical losses and high Kerr nonlinearity, Si3N4 has become one of the main material platforms in hybrid integrated lasers with narrow linewidth [5] and soliton microcombs with low pump threshold [49]. While modulators integrated with Si3N4, e.g. based on the electro-absorption of graphene [212, 73, 72, 214] and ferroelectric PZT [77] have been demonstrated, they have not shown compatibility with soliton generation which requires maintaining ultralow optical losses. As a result, current integrated actuation techniques for soliton microcombs rely primarily on metallic heaters [65, 64], which have limited actuation bandwidth up to 10 kHz. Furthermore, heaters exhibit unidirectional tuning and typically consume electrical power exceeding 30 mW [39], higher than the optical power threshold for soliton formation in state-of-the-art integrated devices [39, 40], and are not compatible with cryogenic operation [67, 254]. Alternatively, the direct generation of solitons in Pockels materials such as AlN [227, 46, 47, 255] and LiNbO3 [45, 44, 43] can allow high-speed actuation, however these platforms are not yet as mature as Si3N4.
@@ 4.1 Narrow-Linewidth, Frequency-Agile Integrated Laser
Over the past decades, the heterogeneous integrated semiconductor laser has revolutionized the application of laser in our daily life, from a laser pointer to a data center, due to its compact size and low cost [256]. However, its frequency noise is still higher than the fiber laser. It is only most recently that the coupling to a low-loss Si3N4 photonic circuit has made it possible for a chip-scale, low-noise laser, thanks to the maturity of hybrid and heterogeneous integration [257]. A variety of coupling schemes has been invented, such as wafer bonding [7, 4, 258, 259], edge coupling [41, 5, 260], and photonic wire bonding [261, 262, 263]. A Hertz-linewidth laser has been achieved by properly tuning the laser injection current such that the CW laser’s frequency coincides with the resonance of an external Si3N4 microresonator [5, 264]. The laser emission frequency is then self-injection-locked to the Si3N4 microresonator, a phenomenon in which the laser dynamics is impacted by external feedback light in the case without an optical isolator [265]. When light is coupled into the microresonator, bulk and surface Rayleigh scattering inside the microresonator can backscatter some portion of the light and inject it into the laser, which triggers laser self-injection locking and impacts the laser dynamics. The consequence is that the new laser emission frequency is tightly locked to the microresonator resonance whose frequency is close to the initial free-running laser’s emission frequency. The laser linewidth and noise are significantly reduced in the presence of this passive optical feedback (in contrast to active locking of laser frequency such as the Pound–Drever–Hall (PDH) lock), and the linewidth is ultimately limited by the thermal-refractive noise (TRN) of the Si3N4 microresonator [266].
@@ 4.1.1 Piezoelectric Tunable Self-Injection Locked DFB Laser
Figure 23(d) shows the laser frequency tuning under a triangularly modulated voltage (150 V) applied on the AlN actuator. The time-varying laser frequency is generally characterised by measuring a heterodyne beatnote with a reference ECDL on a fast photodetector, which produces a time-frequency spectrogram. Consequently, the laser output frequency is triangularly chirped with frequency excursion of 1.1 GHz and speed up to 0.8 MHz. Due to the linearity of the AlN actuator, the frequency chirping nonlinearity is below 10 MHz (defined as the Root Mean Square (RMS) deviation of the measured frequency from an ideal triangular wave calculated by least-squares fitting), and no tuning hysteresis is observed during continuous operation. The degradation of the deviation is primarily due to the nonlinearity of the voltage amplifier at high frequency. This highly linear tuning behavior facilitates the generation of chirps without the need for active or passive linearization, which is ideal for direct implementation in long-range FMCW LiDAR systems [269, 270, 38]. As discussed in the previous chapter, the modulation speed is limited by the chip’s mechanical modes, such as the contour modes [271] and HBAR modes [12]. In Ref. [52], a flat modulation response is extended to 10 MHz by implementing a differential drive on an apodized chip with carbon tape. The linear frequency tuning speed of 1.6 PHz/s is achieved. The tuning speed of the hybrid laser is ultimately limited by the photon lifetime of the cavity and the carrier lifetime of the laser and EHz/s tuning speeds should be possible similar to recent work in electro-optical tuning [217].
In addition to AlN, other piezoelectric materials with larger piezoelectric coefficient, such as lead zirconium titanate (PZT), can be used to reduce the required voltage for a given frequency excursion. The same PZT actuator shown in Fig.19 was used by the same group to tune the frequency of a SIL DFB laser, as demonstrated in Fig. 24(a,b). A 525 MHz frequency excursion at the repetition rate of 500 kHz is achieved by applying only 1.1 V triangular wave (see Fig. 24(b)). The laser tuning was linearized using an iterative algorithm to correct for the intrinsic nonlinearity and hysteresis of the PZT actuator [52]. As a result, the measured RMS nonlinearity is 5 MHz, and the relative nonlinearity, defined as the ratio of the nonlinearity to the total frequency excursion, is less than 1. The small sweeping voltage is also compatible with CMOS integrated circuits (e.g., application-specific integrated circuit (ASIC)), paving the way for a fully photonic-electronic integrated tunable hybrid laser systems [273].